I need help graphings this in Excel, I am not getting the right best fit line equation.

It needs to be cell voltage (Ecell) as a function of Log[Cu+2]

Ecell Log(Cu+2)
1.034 0
1.204 -2
1.017 -3
1.1 -4
1.011 -5
0.992 x
Answer #1

Ecell Vs Log[Cu2+] -6 -5 -4 2

First, plot the given points in Excel. Then right click on any one of the points and select the add trendline option. Automatically a linear trendline would be added.

Identify the encircled points (in black) in the graph, to find the equation of the line y = mx + C

The encircled points (x,y) are (-5, 1.06) and (0,1.09)

Solving for m and C,

For (0, 1.09) , 1.09 = m(0) + C => C = 1.09

For (-5, 1.06), 1.06 = m(-5) + 1.09 => m = 0.006

Thus, the equation of the line is y = 0.006x + 1.09

Solve, this equation for y = 0.992

0.992 = 0.006(x) + 1.09 => x = -16.34